# Just-In-Time Objectives: A General Approach for Specialized AI Interactions

### v1

@@ 4.1.2. Study 2: Exploring Generalizability
In the second evaluation, we conduct a study with an expanded set of online participants (N=205N={205}{}) who provide screenshots of their workspace and evaluate resulting system-generated tools. The online study enrolls participants recruited from Prolific. At the start of the study, participants are asked to upload two screenshots of their desktop that capture a moment in time when they might benefit from AI assistance. These inputs are submitted to our system to perform live objective induction. Then, participants are asked to rate the usefulness and accuracy of the highest-weighted induced objective for their two task screenshots along with the most important goal for this task. Here, they can select from the three top-weighted induced objectives (presented in randomized order) or enter their own custom goal in a text field. The system then generates experts, tool designs, and feedback for the tasks based on the selected goal for each task. The study then takes the same form as the prior study: participants complete a Generator task and an Evaluator task based on their uploaded screenshots. Participants were compensated through Prolific at a \$16/hour rate. The study produced $n=410$ participant submissions, which spanned an even broader range of interfaces (e.g., word processors, spreadsheet software, presentation software, LLM chats, image/audio/video editing tools, design tools, code editors) and task types (e.g., creative writing, academic writing, travel planning, financial planning, health and fitness planning, resume drafting, letter writing).
@@ 4.2. Results
Across the two evaluation settings, we find that our induced goals are rated as highly useful and accurate, and outputs produced with the aid of our just-in-time objective (experts, tool designs, and feedback) perform strongly above those of a baseline unsteered LLM (Figure 8 and Figure 9).
@@ 4.2.1. Just-in-time Objective Accuracy and Usefulness
For Study 2, we have additional insight into which goal users deemed most important for their tasks. We find that just-in-time objectives were chosen in the vast majority of cases (97.8%) over a custom-authored objective (Figure 10). In addition, participants’ selections are well-aligned with the inferred weights associated with induced objectives: when chosen from a randomized ordering, the highest-weighted objective (Obj 1) was selected as the most important in 42.9% of cases, followed by Obj 2 in 29.5% and Obj 3 in 25.4%.
@@ 4.2.3. Just-in-time Objectives for Evaluation
Our best-of-N evaluations explore whether JIT objectives can serve as effective evaluators. In both studies, we find that JIT objectives allow us to sample effectively to produce preferred candidates even from among a pool of JIT generations: Best-of-100 outputs achieve a win rate of 75.0% in Study 1 and 57.6% in Study 2 over Best-of-10 and Best-of-1 (i.e., no evaluation) outputs (Figure 8D, Figure 9D). Across head-to-head comparisons, the higher-N candidate won at higher rates (Study 1: 61.9%, Study 2: 59.2%), indicating that the JIT evaluator helps to sift through the noise and select more user-aligned candidates. However, there may be diminishing returns for additional sampling: in Study 2, Best-of-100 only achieved a 3.5% improvement in win rate. One possibility is that it is already possible to max out performance at a relatively small sample size, so additional samples merely reproduce the same caliber of results. Another possibility is that to better support test-time scaling using the JIT evaluator, we may need to introduce greater variation in our JIT generation process.
Finally, though this study was primarily designed as a rating task, participants proactively shared their excitement about the JIT outputs in an optional final text field of the study, perhaps compelled because the system surfaced AI feedback on tasks of personal importance. Participants indicated that the system surfaced valuable ideas that hadn’t occurred to them before, such as P27: “These were a series of short stories and it never occurred to me to link them all together in one story with a cohesive narrative, I’m thrilled about it!” and P56: “The feedback for the code is fantastic. I didn’t realize that I had vulnerabilities that could be exploited.” Some participants indicated surprise at what the system could do with a screenshot alone: “I’m very impressed at how this AI bot understands what I need to do” (P41), “I’m surprised how helpful these responses are, honestly” (P123). These participant responses present promising evidence that just-in-time objectives can unlock utility to users beyond what they typically encounter with available LLMs.
Figure 10. Participants’ selected objectives are aligned with just-in-time objective inferred weights (N=205N={205}{}). Objectives 1-3 were generated with a JIT objective, where Obj 1 is the highest-weighted candidate, followed by Obj 2, then Obj 3. Objective options were presented in randomized order, and participants could enter a Custom objective if none of the induced options were accurate.:
@@ 4.3. Error Analysis & Limitations
Investigating the generated outputs, we find that experts generated using a JIT objective offered greater task-specific depth than baseline experts. Tool designs generated with JIT objectives succeeded because they tended to align with user objectives rather than document-centric goals. However, Poppins’s specificity could be a double-edged sword: when users were in the early stages of work or performing utilitarian tasks, they often favored baseline tool designs. In such cases, users preferred lower-level, low-effort aids (e.g., a “character voice generator”) that the baseline tended to produce rather than higher-level, meta-cognitive tools (e.g., a “character development framework”) that a JIT objective would often produce. This trend was most pronounced in brainstorming and text comprehension, where users preferred utilitarian baseline tools. Feedback, by contrast, showed the most variable performance: we found no consistent relationship between task type or domain and preference, suggesting that feedback preferences may be more user-dependent.
@@ 5. EVALUATION: Powering UI Generation with Just-In-Time Objectives
Our in-lab sessions explore whether just-in-time objectives allow our Poppins system to produce tools that are useful and relevant to individual users.
@@ 5.1. Procedure
We conduct 1-hour sessions with participants (N=17N={17}{}) where they apply provided tools to their own writing task. Participants experiment with both using these generated tools and modifying them to refine their behavior.
@@ 5.1.1. Protocol
Our recruitment process is described in Section 4.1.1. The main study consists of a 1-hour session with each participant to evaluate system-generated tools and baseline outputs:
@@
2: Baseline Comparison phase (30 min). The first phase compares tools that incorporate just-in-time objectives (Poppins) with those that do not (Baseline). Condition order is randomized across participants (Baseline-Poppins-Interview or Poppins-Baseline-Interview).
Figure 11. From participant-provided writing tasks, Poppins-tools generated a wide range of tools, including a Character Emotion Tracker, Neural Architecture Search Method Explorer, and Character Voice Consistency Checker. These tools were not only highly varied and user-specific, but directly useful for participants’ work.:
@@
    • Create and review assistance: Participant requests assistance from the baseline tool and reviews the output.
    • Iterate on assistance: Participant modifies the assistance to better align with their needs.
    • Rate tool: Participant completes a brief survey rating their experience with this tool.
  • Poppins-experts. In this phase, the participant uses the Poppins-experts mode, which performs objective induction and goal-based generation and evaluation to generate expertise for assistance. This system mode selects from a pre-implemented set of UI design templates (including Feedback, Brainstorming, and Line Editor interfaces), which is useful to provide consistency and facilitate comparison across participants.
@@
    • Create and review tool: Participant creates a tool from their writing task context and reviews the output.
    • Iterate on tool: Participant modifies the tool to better align with their needs.
    • Rate tool: Participant completes a brief survey rating their experience with this tool.
  • Interview. We conclude this phase with an interview to understand participants’ impressions of both tools.
3: Exploratory phase (20 min). We conduct another task with the Poppins-tools mode as a design probe. For this phase, our goal is to explore whether just-in-time objectives can perform the full UI generation task without intervention and how participants respond to this design concept.
@@
    • Create and review tool: Participant creates a tool from their writing task context and reviews the output.
    • Rate tool: Participant completes a brief survey rating their experience with this tool.
  • Interview. We conclude with an interview to understand participants’ holistic experience with Poppins and the baseline, as well as gather feedback on how Poppins may fit into their everyday computer use.
Figure 12. Both Poppins variants produce highly relevant goals, tool designs, and model expertise. Participants find Poppins-experts highly relevant and substantially more useful than Poppins-tools, even with the same underlying just-in-time objectives. These results suggest that interface quality (more inconsistent with Poppins-tools) may strongly influence assessments).:
@@ 5.2. Results
To address whether systems that incorporate just-in-time objectives are useful and relevant to users’ needs, we analyzed both the goals generated and the resulting system outputs. Our findings indicate that Poppins produces marked shifts in user experience by surfacing latent intentions and reducing prompting burden, all the while producing more detailed, specific, and interpretable outputs.
Participants brought in a range of different writing tasks to the session, such as a short story, draft research paper, scholarship application, presentation script, reading notes, and manga translations. With Poppins, these tasks produced a broad range of different tools for assistance, for example: an Academic Jargon Organizer to analyze a presentation script and improve language accessibility, a Neural Architecture Search (NAS) Explorer to explore different NAS methods alongside expert insights, a Concept Visualizer to transform abstract theoretical concepts into concrete visual diagrams, and a Character Voice Consistency Checker to analyze character dialogue patterns to maintain consistent personality traits (Figure 11).
@@ 5.2.1. Are Poppins outputs relevant?
Overall, we find that the system’s induced goals are highly relevant to user goals. By achieving accurate goals and rendering them as functional levers, Poppins supports more fruitful and cognitively aligned interactions with an LLM. Across both the Poppins-experts (P-E) and Poppins-tools (P-T) conditions, all participants rated the system’s induced goals as at least “somewhat relevant,” with the vast majority rating the induced goals as at least“relevant” (P-E: 94.1%, P-T: 82.4%) or “very relevant” (P-E: 64.7%, P-T: 52.9%) (Figure 12). We observe similar trends for the proposed tool and expertise, with the vast majority of participants rating these as “relevant” or “very relevant” (P-E tools: 82.4%, P-T tools: 88.2%; P-E expertise: 76.5%, P-T expertise: 82.4%).
Induced goals capture nascent user intentions
Poppins proved highly effective at surfacing objectives that align well with user goals. Most notably, participants reported that Poppins articulated important but subconscious objectives they would struggle to express explicitly: “Initially, I didn’t have any goals, but [Poppins] actually gave me goals where I felt like oh, some of it actually aligns with what I want […] It actually is something that I need, but I just didn’t know until I saw it” (P18). This ability to express the inexpressible seems particularly promising for participants who felt this was a fundamental constraint on their usage of LLMs, such as P10 who shared, “I only use large language models for tasks where I can clearly explain what I want to do.” In contrast to traditional text prompts that capture only a narrow slice of user intent, Poppins would sometimes “read the user’s mind” to surface useful and surprisingly astute objectives. Objectives provided a helpful launching-off point when users felt uncertain on where they ought to focus their attention or seek assistance: “Sometimes I feel like I don’t even know what kind of stuff I want to change or what kind of direction I am going, so if they do the first step for me (infer my goals for this writing), it’s definitely a good starting point” (P5).
Goal-based levers shift users from thinking about prompting to thinking about their task
Because induced goals were strongly aligned with users’ own goals, Poppins enabled users to redirect cognitive effort from low-level prompt engineering toward higher-level objective refinement. As P25 explained, they were able to “focus on what kind of feedback [they] wanted…rather than how [they] needed to word [their] prompt to the LLM to get the feedback.” Participants sometimes found this shift surprising or counterintuitive, as it felt like they could invest less effort with Poppins, but get out higher-quality results:
@@
Figure 13. Poppins-experts produces outputs that are higher quality, more useful, and yield greater control than a baseline LLM. Poppins-tools maintains higher control than the baseline, but inconsistent interface quality may detract from its usefulness and relevance.:

### v2

@@ 4.1.2. Study 2: Exploring Generalizability
In the second evaluation, we conduct a study with an expanded set of online participants (N=205N={205}{}) who provide screenshots of their workspace and evaluate resulting system-generated outputs. The online study enrolls participants recruited from Prolific. At the start of the study, participants are asked to upload two screenshots of their desktop that capture a moment in time when they might benefit from AI assistance. These inputs are submitted to our system to perform live objective induction. Then, participants are asked to rate the usefulness and accuracy of the highest-weighted induced objective for their two task screenshots along with the most important goal for this task. Here, they can select from the three top-weighted induced objectives (presented in randomized order) or enter their own custom goal in a text field. The system then generates experts, tool designs, and feedback for the tasks based on the selected goal for each task. The study then takes the same form as the prior study: participants complete a Generator task and an Evaluator task based on their uploaded screenshots. Participants were compensated through Prolific at a \$16/hour rate. The study produced $n=410$ participant submissions, which spanned an even broader range of interfaces (e.g., word processors, spreadsheet software, presentation software, LLM chats, image/audio/video editing tools, design tools, code editors) and task types (e.g., creative writing, academic writing, travel planning, financial planning, health and fitness planning, resume drafting, letter writing).
@@ 4.2. Results
Across the two evaluation settings, we find that our induced objectives are rated as highly useful and accurate, and outputs produced with the aid of our just-in-time objective (experts, tool designs, and feedback) perform strongly above those of a baseline unsteered LLM (Figure 8 and Figure 9).
@@ 4.2.1. Just-in-time Objective Accuracy and Usefulness
For Study 2, we have additional insight into which goal users deemed most important for their tasks. We find that just-in-time objectives were chosen in the vast majority of cases (97.8%) over a custom-authored objective (Figure 10). These results cannot be interpreted as 97.8% accuracy, as authoring a custom objective is more effortful than selecting an existing objective. However, this finding indicates that induced objectives were sufficiently relevant such that almost all participants did not opt to write their own objective. In addition, participants’ selections are well-aligned with the inferred weights associated with induced objectives: when chosen from a randomized ordering, the highest-weighted objective (Obj 1) was selected as the most important in 42.9% of cases, followed by Obj 2 in 29.5% and Obj 3 in 25.4%.
Figure 10. Participants’ selected objectives are aligned with just-in-time objective inferred weights (N=205N={205}{}). Objectives 1-3 were generated with a JIT objective, where Obj 1 is the highest-weighted candidate, followed by Obj 2, then Obj 3. Objective options were presented in randomized order, and participants could enter a Custom objective if desired.
Participant objective selection aligned with system-inferred weights across 205 participants. A bar chart showing participant selection frequency for different objective options. Objective 1 (highest weighted) was selected 43% of the time, Objective 2 selected 30%, Objective 3 selected 25%, and custom objectives only 2%.
@@ 4.2.2. Just-in-time Objectives for Generation
Qualitatively, the baseline condition produces results that are relevant to a user’s context, but JIT objectives produce results that advance a user’s specific goals. For example, one participant submitted a screenshot of TA training handbook draft that had an unresolved comment about clarifying the role of TAs. The baseline produced outputs that are relevant to a TA handbook, but do not address this action item: a tool for helping TAs understand different course options, feedback on teaching strategies to add to the document, and expertise from a first-time TA. However, because the induced JIT objective was “Enhance TA role description,” the JIT generator produced outputs that were much more useful to the participant: a tool to expand descriptions of TA responsibilities with concrete examples of professional behavior, feedback on specific examples that could be added to the TA role, and expertise from an experienced TA.
@@ 4.2.3. Just-in-time Objectives for Evaluation
Our best-of-N evaluations explore whether JIT objectives can serve as effective evaluators. In both studies, we find that JIT objectives allow us to sample effectively to produce preferred candidates even from among a pool of JIT generations: Best-of-100 outputs achieve a win rate of 75.0% in Study 1 and 57.6% in Study 2 over Best-of-10 and Best-of-1 (i.e., no evaluation) outputs (Figure 8D, Figure 9D). Across head-to-head comparisons, the higher-N candidate won at higher rates (Study 1: 61.9%, Study 2: 59.2%), indicating that the JIT evaluator helps to sift through the noise and select more user-aligned candidates. However, there may be diminishing returns for additional sampling: in Study 2, Best-of-100 only achieved a 3.5% improvement in win rate. One reason may be that it is already possible to max out performance at a small sample size for these tasks, so additional samples merely reproduce the same caliber of results. Another possibility is that to better support test-time scaling using the JIT evaluator, we may need to introduce greater variation in our JIT generations.
Finally, though this study was primarily designed as a rating task, participants proactively shared their excitement about the JIT outputs in an optional final text field of the study, perhaps compelled because the system surfaced AI feedback on tasks from their own daily lives. Participants indicated that the system surfaced valuable ideas that hadn’t occurred to them before, such as P27: “These were a series of short stories and it never occurred to me to link them all together in one story with a cohesive narrative, I’m thrilled about it!” and P56: “The feedback for the code is fantastic. I didn’t realize that I had vulnerabilities that could be exploited.” Some participants indicated surprise at what the system could do with a screenshot alone: “I’m very impressed at how this AI bot understands what I need to do” (P41), “I’m surprised how helpful these responses are, honestly” (P123). These participant responses present promising evidence that just-in-time objectives can unlock utility to users beyond what they typically encounter with available LLMs.
@@ 4.3. Error Analysis & Limitations
Investigating the generated outputs, we find that experts generated using a JIT objective offered greater task-specific depth than baseline experts. Tool designs generated with JIT objectives succeeded because they tended to facilitate a user’s specific goals rather than context-related actions. However, this specificity could be a double-edged sword: when users were in the early stages of work (e.g., brainstorming) or performing straightforward tasks (e.g., text comprehension), they often favored baseline tool designs. In such cases, users preferred lower-level, low-effort aids (e.g., “character voice generator”) that the baseline tended to produce rather than higher-level, meta-cognitive tools (e.g., “character development framework”) that a JIT objective would often produce. Feedback, by contrast, showed the most variable performance: we found no consistent relationship between task type or domain and preference, suggesting that feedback preferences may be more user-dependent.
Our JIT generator evaluation (Section 4.2.2) runs a comparison to assess whether JIT generators improve upon baseline LLMs. Meanwhile, the JIT evaluator evaluation (Section 4.2.3) is designed to assess whether there is additional benefit in running a JIT evaluator on top of a JIT generator to verify that outputs match an induced objective. This evaluation design does not compare a JIT evaluator against a generic LLM-as-a-judge evaluator (i.e., one that selects “the best candidate” with no knowledge of the JIT objective). However, we qualitatively observe that JIT-selected outputs were successful by strongly encoding the objective, which was not necessarily aligned with a generic notion of quality.
To further investigate this difference, we additionally ran a follow-up analysis on Study 2 using a generic LLM-as-a-judge prompt without the JIT objective. In cases where the higher-N result was preferred (in other words, where the user validated that the JIT evaluator had selected a better candidate), a performant evaluator should achieve strong agreement. We find that the generic evaluator prompt achieved only moderate agreement with the JIT evaluator on these cases ($\kappa$ = 0.46) and overall ($\kappa$ = 0.34), indicating that the generic evaluator displays different selection behavior and would not perform equivalently to the JIT evaluator.
@@ 5. EVALUATION: Powering UI Generation with Just-In-Time Objectives
Our in-lab sessions explore whether JIT objectives allow our Poppins system to produce tools that are useful and relevant to users.
@@ 5.1. Procedure
We conduct 1-hour sessions with participants (N=17N={17}{}) where they apply provided tools to their own writing task. Participants experiment with both using these generated tools and modifying them to refine their behavior. This study was designed for qualitative feedback and was not intended or powered for statistical comparisons across conditions.
@@ 5.1.1. Protocol
Our recruitment process is described in Section 4.1.1. We evaluate system-generated and baseline outputs in a 1-hour session with each participant according to the following structure:
@@
2: Baseline Comparison Phase (30 min). The first phase compares tools that incorporate just-in-time objectives (Poppins) with those that do not (Baseline). Condition order in this comparison phase is randomized across participants (Baseline-Poppins-Interview or Poppins-Baseline-Interview) as a counterbalancing measure for potential learning effects.
@@
    • Generate and review: Participant requests assistance from the baseline tool and reviews the output.
    • Iterate: Participant iterates on chat responses to better align LLM outputs with their needs.
    • Rate: Participant completes a brief survey rating their experience with this system variant.
  • Poppins-experts. In this phase, the participant uses Poppins-experts, which performs objective induction and JIT generation and evaluation to generate expert responses. As described in Section 3.2.2, Poppins-experts selects from a pre-implemented set of output formats, which provides greater consistency to facilitate comparison across participants.
@@
    • Generate and review: Participant requests expert responses for their writing task context and reviews the output.
    • Iterate: Participant iterates on expert generation with Poppins-experts to better align expert responses with their needs. Here, participants could modify any intermediate representations in the system (objectives, experts, output formats) to re-generate expert responses.
    • Rate: Participant completes a brief survey rating their experience with this system variant.
  • Interview. We conclude this phase with an interview to understand the participant’s impressions of both conditions.
3: Exploratory Phase (20 min). We conduct another task with Poppins-tools as a design probe. For this phase, we explore whether JIT objectives can perform the full UI generation task without intervention and how participants respond to this design concept.
@@
    • Generate and review: Participant requests a tool for their writing task context and reviews the output.
    • Rate: Participant completes a brief survey rating their experience with this system variant.
  • Interview. We conclude with an interview to understand the participant’s holistic experience with Poppins and the baseline, as well as gather feedback on how Poppins may fit into their everyday computer use.
@@ 5.2. Results
Figure 11. From participant-provided writing tasks, Poppins-tools generated a wide range of tools, including a Character Emotion Tracker, Neural Architecture Search Method Explorer, and Character Voice Consistency Checker. These tools were not only highly varied and user-specific, but directly useful for participants’ work.
Examples of diverse tools generated by Poppins-tools for participant writing tasks. Three screenshots showing different generated tools: Character Emotion Tracker for analyzing emotional arcs in creative writing, Neural Architecture Search Method Explorer for research projects, and Character Voice Consistency Checker for maintaining dialogue patterns. Each tool displays specialized interface elements tailored to the specific writing task and user context.
To address whether systems that incorporate just-in-time objectives are useful and relevant to users’ needs, we analyzed both the induced objectives and the resulting system outputs. Our findings indicate that Poppins produces marked shifts in user experience by inducing useful objectives and reducing prompting burden, all the while producing more detailed, specific, and interpretable outputs.
Participants brought in a range of different writing tasks to the session, such as a short story, draft research paper, scholarship application, presentation script, reading notes, and manga translations. With Poppins, these tasks produced a broad range of different tools for assistance, for example: an Academic Jargon Organizer to analyze a presentation script and improve language accessibility, a Neural Architecture Search (NAS) Explorer to explore different NAS methods alongside expert insights, a Concept Visualizer to transform abstract theoretical concepts into concrete visual diagrams, and a Character Voice Consistency Checker to analyze character dialogue patterns to maintain consistent personality traits (tool examples in Figures 11, 15).
@@ 5.2.1. Are Poppins outputs relevant?
Overall, we find that the system’s induced objectives are highly relevant to user goals. By achieving accurate objectives and rendering them as functional levers, Poppins supports more fruitful and cognitively aligned interactions with an LLM. For both Poppins-experts (P-E) and Poppins-tools (P-T), all participants rated the system’s induced objectives as at least “somewhat relevant,” with the vast majority rating the induced objectives as at least“relevant” (P-E: 94.1%, P-T: 82.4%) or “very relevant” (P-E: 64.7%, P-T: 52.9%) (Figure 12). We observe similar trends for the proposed tool and expertise, with the vast majority of participants rating these as “relevant” or “very relevant” (P-E tools: 82.4%, P-T tools: 88.2%; P-E expertise: 76.5%, P-T expertise: 82.4%).
Induced objectives lower the barrier to specifying intents
Poppins proved highly effective at surfacing induced objectives that align well with user objectives. Most notably, participants reported that Poppins articulated important objectives they would struggle to express explicitly: “Initially, I didn’t have any goals, but [Poppins] actually gave me goals where I felt like oh, some of it actually aligns with what I want […] It actually is something that I need, but I just didn’t know until I saw it” (P18). Lowering the effort barrier to intent specification seems particularly promising for participants who felt this was a fundamental constraint on their usage of LLMs, such as P10 who shared, “I only use large language models for tasks where I can clearly explain what I want to do.” In contrast to traditional text prompts that capture a narrow slice of user intent with active user effort, Poppins automatically surfaced useful objectives rich in user-specific detail. Objectives provided a helpful launching-off point when users felt uncertain on where they ought to focus their attention or seek assistance: “Sometimes I feel like I don’t even know what kind of stuff I want to change or what kind of direction I am going, so if they do the first step for me (infer my goals for this writing), it’s definitely a good starting point” (P5).
Figure 12. Both Poppins variants produce objectives, tool designs, and model expertise that are relevant and useful.
Relevance and Usefulness ratings for Poppins-generated goals, tool designs, and expertise across both system variants. Two sets of stacked bar charts comparing Poppins-experts and Poppins-tools across three output types: objectives, tools, and expertise. Both systems show high relevance and usefulness ratings, with a majority of ”Agree” and ”Strongly agree” ratings from participants.
Objective-based levers shift users from thinking about prompting to thinking about their task
Because induced objectives were strongly aligned with users’ goals, Poppins enabled users to redirect cognitive effort from low-level prompt engineering toward higher-level objective refinement. As P25 explained, they were able to “focus on what kind of feedback [they] wanted…rather than how [they] needed to word [their] prompt to the LLM to get the feedback.” Participants sometimes found this shift surprising, as it felt like with Poppins, they could invest less effort, but receive higher-quality results:
@@
Figure 13. Poppins produces high-quality, useful, controllable, and relevant assistance to participants.
User experience comparison between Poppins variants and Baseline system across overall quality, usefulness, user control, and relevance. Sets of three stacked bar charts comparing Baseline, Poppins-experts, and Poppins-tools for all four measures. Poppins-experts outperforms Baseline and Poppins-tools on Overall Quality and Usefulness, while both Poppins variants provide greater User Control than the baseline system.
Figure 14. Comparing participant Poppins-experts vs. Baseline survey responses (Poppins - Baseline), Poppins produces higher overall quality, user control, usefulness, and relevance ratings than the baseline, which already performed quite strongly.
Changes in ratings between Poppins-experts and Baseline conditions for overall quality, user control, usefulness, and relevance. Each plot displays the median, an 95% bootstrapped confidence interval, and datapoints for individual participants. All metrics display a positive change in ratings for Poppins-Experts compared to Baseline.